OpenStudy (anonymous):

for the line segment whose endpoints are R(1,2) and S(6,7), find the y value for the point located 3/4 the distance from R to S. a) 4.25 b)4.75 c)5.25 d)5.75

OpenStudy (anonymous):

@hero

OpenStudy (jdoe0001):

same as before http://openstudy.com/updates/539f5d85e4b0eb8de56e97e0

OpenStudy (anonymous):

would the ratio be 3:4 ?

OpenStudy (jdoe0001):

the line is again, split in 4 equal pieces say to point P then from RP 3/4 and 1/4 for PS

OpenStudy (jdoe0001):

so you'd do it the same way as the previous one

OpenStudy (jdoe0001):

just change the points values and the ratio now is not 1:3 BUT D 3:1 because is from R to S

OpenStudy (anonymous):

ooo ok that makes sense

OpenStudy (anonymous):

ok so i got (2.25, 3.25).. i dont think it is right @jdoe0001

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

\(\bf R(1,2)\qquad S(6,7)\qquad ratio1=3\qquad ratio2=1\qquad 3:1\\ \quad \\ \quad \\ \cfrac{RP}{SP}=\cfrac{ratio1}{ratio2}\implies ratio2\cdot RP=ratio1\cdot SP\quad \textit{dividing by P}\\ \quad \\ ratio2\cdot R=ratio1\cdot S\implies 1(1,2)=3(6,7)\\ \quad \\\qquad {\color{blue}{ P=\left(\cfrac{\textit{sum of "x" values}}{ratio1+ratio2}\quad ,\quad \cfrac{\textit{sum of "y" values}}{ratio1+ratio2}\right)}}\\ \quad \\ \qquad thus\qquad \\ \quad \\ P=\left(\cfrac{(1\cdot 1)+(3\cdot 6)}{3+1}\quad ,\quad \cfrac{(1\cdot 2)+(3\cdot 7)}{3+1}\right)\)

OpenStudy (jdoe0001):

recalculate your y-coordinate, see what you get

OpenStudy (anonymous):

5.75!! thank you!

OpenStudy (jdoe0001):

yw